# Neutrinoless $\beta\beta$ decay nuclear matrix elements complete up to N2LO in heavy nuclei

### v1

@@ 4.3 Combined N2LO NMEs
With the results in Table 2 we can compare the relative size of the $0\nu\beta\beta$ NMEs. In particular, the largest N2LO contributions are typically the ultrasoft NME for the NSM, and the CT part of the loops NME for the pnQRPA. They amount to $\lesssim 10\%$ and $\lesssim 20\%$ of the standard long-range LO NME, respectively, which is somewhat larger than expected from the $\chi$EFT power counting [35, 41]. The ultrasoft N2LO NME is especially large for 100Mo—about 50% of the LO long-range NME. However, even in this extreme case the combined N2LO NME is about a third of the total LO NME.
The size of the largest N2LO NMEs is nonetheless similar to the impact of the nucleon dipole form factors used for $M^{0\nu}_{\rm long}$, typically a $10\%-15\%$ effect [54, 68], from diagrams that enter mainly at the same order. The magnetic moment operator, a N2LO term as well, also contributes at the $10\%$ level for several nuclei [44, 46]. In addition, in Sec. 4.1 we discussed that, for the NSM, the relatively large ultrasoft NMEs are in good agreement with the expectation of beyond closure calculations.
Figure 4: Relative contributions of the ultrasoft (red and dark blue bars) and complete N2LO (orange and light blue) NMEs compared to the LO NME, obtained with the NSM and QRPA for all nuclei studied in this work. The bars show the uncertainties in Table 1.
Figure 4 summarizes for the ratios of the N2LO and total LO NMEs, with the uncertainty ranges from Table 1. The relative impact of $M^{0\nu}_{\rm usoft}$ is typically a $5\%-10\%$ enhancement (pnQRPA, blue bars) or reduction (NSM, red bars). Likewise, the central values of the combined N2LO contributions also differ in sign, even though the NSM errorbars (orange) allow for a positive contribution as predicted by the pnQRPA (light blue). The central values of the total N2LO NMEs are about $-5\%$ for the NSM and $+10\%$ for the pnQRPA, with the extreme absolute values being limited to $\leq 20\%$ and $\leq 25\%$, respectively, except in the extreme case of 100Mo where it can reach up to $50\%$ of the LO NME.
@@ 5 Summary and outlook
We have calculated for the first time complete $0\nu\beta\beta$ NMEs up to N2LO for nine medium-mass and heavy nuclei including the most promising experimental candidates 76Ge, 100Mo, 130Te and 136Xe. Our NSM results support the $\chi$EFT prediction that the NMEs from ultrasoft neutrinos effectively capture most contributions beyond the closure approximation, while in the pnQRPA $M^{0\nu}_{\rm usoft}$ also has the same sign as the beyond-closure effects, but overestimates them by about a factor two.
The absolute central value of the ultrasoft NMEs ($5\%-10\%$ of the total LO NME) are typically larger than the NMEs coming from loop diagrams that cannot be absorbed into nuclear form factors. However, since the latter are rather short-ranged, they are associated with sizeable uncertainties due to the regulator value and SRCs, and the extreme absolute values of $M^{0\nu}_{\rm usoft}$ and $M^{0\nu}_{\rm loops}$ are comparable.
The main discrepancy between the NSM and pnQRPA N2LO NMEs is that for the former $M^{0\nu}_{\rm usoft}$ has opposite sign to the LO NMEs, while for the latter it has the same sign. This is due to the different behaviour of the states in the intermediate $0\nu\beta\beta$ nucleus, which could be probed experimentally in charge-exchange reactions. The difference between the results of the two many-body methods persists for the combined N2LO NMEs, which typically range between $-15\%$ to $+5\%$ in the NSM, and $0\%$ to $+20\%$ for the pnQRPA. Our N2LO NMEs are anomalously large for 100Mo: the ultrasoft term can be up to 30% and the combined N2LO contribution up to 50%. Overall, N2LO NMEs should be included in any $0\nu\beta\beta$ calculation since their impact is moderate but comparable to the one of other N2LO diagrams, like nucleon form factors, routinely included in NME studies.
For the future, we would like to narrow down the uncertainties in $M^{0\nu}_{\rm loops}$ by investigating more carefully the dependence with different regulators and regulator forms. Additionally, we would like to include two-body currents, which in the $\chi$EFT power counting enter at N3LO [35, 42]. Nonetheless, in $\beta$ decays the importance of two-body currents is critical to understand the overestimation of matrix elements in many-body calculations (or “quenching”) [75, 76]. In fact, based on the impact on $\beta$ decays, some estimations for $0\nu\beta\beta$ decay suggest an effect comparable to the N2LO NMEs studied in this work [77, 78, 23], even though smaller contributions are also possible [79].
@@ Acknowledgements
We would like to thank A. Belley, V. Cirigliano, W. Dekens, M. Drissi and J. Engel for insightful discussions. We thank the Institute for Nuclear Theory at the University of Washington for its kind hospitality and stimulating research environment. This research was supported in part by the INT’s U.S. Department of Energy grant No. DE-FG02- 00ER41132. This work was supported by Arthur B. McDonald Canadian Astroparticle Physics Research Institute, the MCIN/AEI/10.13039/5011 00011033 from the following grants: PID2020-118758GB-I00, PID2021-127890NB-I00 and RYC-2017-22781 through the “Ramón y Cajal” program funded by FSE “El FSE invierte en tu futuro”, CNS2022-135716 funded by the “European Union NextGenerationEU/PRTR”, and CEX2019-000918-M to the “Unit of Excellence María de Maeztu 2020-2023” award to the Institute of Cosmos Sciences; and by the Generalitat de Catalunya, grant 2021SGR01095. TRIUMF receives federal funding via a contribution agreement with the National Research Council of Canada.

### v2

@@ 4.3 Combined N2LO NMEs
Figure 4: Relative contributions of the total ultrasoft ($M^{0\nu}_{\rm tot,usoft}$, red and dark blue bars) and complete N2LO (orange and light blue) NMEs compared to the LO NME, obtained with the NSM and pnQRPA for all nuclei studied in this work. The bars show the uncertainties in Table 1.
With the results in Table 2, we can compare the relative size of the $0\nu\beta\beta$ NMEs. In particular, the largest N2LO contributions are typically the ultrasoft NME for the NSM, and the CT part of the loops NME for the pnQRPA. They amount to $\lesssim 10\%$ and $\lesssim 20\%$ of the standard long-range LO NME, respectively, which is somewhat larger than expected from the $\chi$EFT power counting [36, 42]. The ultrasoft N2LO NME is especially large for 100Mo—about 50% of the LO long-range NME. However, even in this extreme case the combined N2LO NME is about a third of the total LO NME.
The size of the largest N2LO NMEs is nonetheless similar to the impact of the nucleon dipole form factors used for $M^{0\nu}_{\rm long}$, typically a $10\%$-$15\%$ effect [55, 67], from diagrams that enter mainly at the same order. The magnetic moment operator, a N2LO term as well, also contributes at the $10\%$ level for several nuclei [45, 47]. In addition, in Sec. 4.1 we discussed that, for the NSM, the relatively large ultrasoft NMEs are of similar size as beyond closure effects.
[…]
@@ 5 Summary and outlook
We have calculated for the first time complete $0\nu\beta\beta$ NMEs up to N2LO for nine medium-mass and heavy nuclei including the most promising experimental candidates 76Ge, 100Mo, 130Te and 136Xe.
The absolute central value of the ultrasoft NMEs ($5\%$-$10\%$ of the total LO NME) are typically larger than the NMEs coming from loop diagrams that cannot be absorbed into nuclear form factors. Both terms introduce sizable uncertainties, due to their ultrasoft-scale dependence or short-range character. When considering theoretical uncertainties, ultrasoft NMEs are of similar size and same sign as beyond-closure effects.
The main discrepancy between the NSM and pnQRPA N2LO NMEs is that for the former $M^{0\nu}_{\rm tot,usoft}$ has opposite sign to the LO NMEs, while for the latter it has the same sign. This is due to the different behaviour of the states in the intermediate $0\nu\beta\beta$ nucleus, which could be probed experimentally in charge-exchange reactions. The difference between the results of the two many-body methods persists for the combined N2LO NMEs, which typically range between $-20\%$ to $+5\%$ in the NSM, and $0\%$ to $+30\%$ for the pnQRPA. Our N2LO NMEs are anomalously large for 100Mo: the ultrasoft term can be up to 40% and the combined N2LO contribution up to 60%. Overall, N2LO NMEs should be included in any $0\nu\beta\beta$ calculation since their impact is moderate but comparable to the one of other N2LO diagrams, like nucleon form factors, routinely included in NME studies.
For the future, we would like to narrow down the uncertainties in $M^{0\nu}_{\rm loop}$ by investigating more carefully the dependence with different regulators and regulator forms. We would also like to use $\chi$EFT Hamiltonians to check that the $\mu_{\text{us}}$ scale dependence of $M^{0\nu}_{\rm tot,usoft}$ vanishes. Additionally, we would like to include two-body currents, which in the $\chi$EFT power counting enter at N3LO [36, 43]. Nonetheless, in $\beta$ decays the importance of two-body currents is critical to understand the overestimation of matrix elements in many-body calculations (or “quenching”) [74, 75]. In fact, based on the impact on $\beta$ decays, some estimations for $0\nu\beta\beta$ decay suggest an effect comparable to the N2LO NMEs studied in this work [76, 77, 23], even though smaller contributions are also possible [78].
@@ Acknowledgements
We would like to thank A. Belley, V. Cirigliano, W. Dekens, M. Drissi and J. Engel for insightful discussions. We thank the Institute for Nuclear Theory at the University of Washington for its kind hospitality and stimulating research environment. This research was supported in part by the INT’s U.S. Department of Energy grant No. DE-FG02- 00ER41132. This work was supported by Arthur B. McDonald Canadian Astroparticle Physics Research Institute, the MCIN/AEI/10.13039/5011 00011033 from the following grants: PID2020-118758GB-I00, PID2023-147112NB-C22, RYC-2017-22781 through the “Ramón y Cajal” program funded by FSE “El FSE invierte en tu futuro”, CNS2022-135716 funded by the “European Union NextGenerationEU/PRTR”, and CEX2019-000918-M to the “Unit of Excellence María de Maeztu 2020-2023” award to the Institute of Cosmos Sciences; and by the Generalitat de Catalunya, grant 2021SGR01095. TRIUMF receives federal funding via a contribution agreement with the National Research Council of Canada.